OpenStudy (anonymous):

Consider right triangle ABC with C = 90°. If a = 6 and b = 8, find c. Please help if you can, Thanks!

OpenStudy (sleepyjess):



OpenStudy (sleepyjess):



OpenStudy (sleepyjess):

Now use Pythagorean Theorem, \(\sf 6^2~+~8^2~=~c^2\\36~+~64~ =~c^2\\100~=~c^2\) Now take the square root of 100 and the square root of \(\sf c^2\) and you have x = c

OpenStudy (anonymous):

Thanks! This actually makes alot more sense now :)

OpenStudy (sleepyjess):

No problem! Welcome to OpenStudy!

OpenStudy (sleepyjess):

Since you got help, could you please close this to let others know you have been helped?